OpenStudy (anonymous):

you commute 56 miles one way to work. the trip takes 12 minutes longer than the trip home. your average speed on the trip home was 6 miles per hour slower. what is your average speed on the trip home

OpenStudy (anonymous):

so the formula is d=rt meaning distance = rate times time

OpenStudy (anonymous):

i got to this far. to work: 56=r(t-12/60) to home: 56=(r-6)t but i dont know how to solve it...